OpenStudy (anonymous):

A map has a scale of 1 cm : 20 km. Two cities are 2.5 cm apart on the map. To the nearest tenth of a kilometer, what is the actual distance corresponding to the map distance?

OpenStudy (smurfy14):

set it up as a porportion 1/20=2.5/x then cross multiply so x=20(2.5) x=50km